Numerical problem on consumer surplus: Assume that the demand for travel over a



bridge takes the form Y = 1,000,000 – 50,000P, where Y is the number of trips over the



bridge and P is the bridge toll (in dollars).




a. Calculate the consumer surplus if the bridge toll is $0, $1, and $20.




b. Assume that the cost of the bridge is $1,800,000. Calculate the toll at which the bridge



owner breaks even. What is the consumer surplus at the breakeven toll?




c. Assume that the cost of the bridge is $8 million. Explain why the bridge should be built



even though there is no toll that will cover the cost

1
Expert's answer
2022-03-03T12:36:19-0500

a)

Y = 1,000,000 - 50,000P


Y = Represents the number of trips over the bridge

 P = Represents the bridge toll


Inverse demand is calculated as;

"-50,000 P = Y - 1,000,000"


"P =20 - \frac{Y}{50,000}"


To find the consumer surplus

"Consumer surplus = ½ \times base \times height"


Base = Y ( number of trips per price)

Height = ( willingness price - Actual price)


We need to get the willingness price when Y =0


"= P = 20 - \frac{Y}{50,000}"


"P = 20 - \frac{0}{50,000}"


"P = 20"

 


Consumer surplus when the toll is $0


"Consumer surplus = ½ \times base \times height"


"Consumer surplus = ½ \times 1,000,000 \times (20-0)"


"Consumer surplus =10,000,000"



Consumer surplus when the toll is $1


"Consumer surplus = ½ \times base \times height"


"Consumer surplus = ½ \times 950,000 \times (20-1)"


"Consumer surplus =9,025,000"



Consumer surplus when the toll is $20

"Consumer surplus = ½ \times base \times height"


"Consumer surplus = ½ \times 0 \times (20-20)"


"Consumer surplus =0"



 b)

At Break even point

"= TR - TC = 0"


"TR = P\times Y"


"TR = 20 - \frac{Y}{50,000}Y"


"TR = 20Y - \frac{Y^2} {50,000}"


"0 = 20Y - \frac{Y^2} { 50,000} - 1,800,000"


"1,800,000 =20Y - \frac{Y^2} {50,000}"


"90,000,000,000 = 1,000,000Y - Y2"


"-Y^2 +1000,000Y - 90,000,000,000"



Use the quadratic formula


"X = -b +- √ \frac {b^2 -4ac} {2a}"


"X = - 1,000,000 +- √\frac { 1,000,0002 -4\times -1\times 90,000,000,000} {2\times -1}"


"X = 100,000 units\space or \space900,000 \space units"


"Y = 900,000 \space or\space 100,000"



We find the price by replacing Y


"P = 20 - \frac{Y}{50,000}"


"P = 20 - \frac {900,000}{50,000}"


"p=2"



Or

"P = 20 - \frac {Y}{50,000} \space

P = 20 - \frac {100,000}{50,000}"


"P = 18"



Then we find the consumer surplus


Consumer surplus when the toll is $2


"Consumer surplus = ½ \times base \times height"


"Consumer surplus = ½ \times 900,000 \times (20-2)"


"Consumer surplus =8,100,000"



Consumer surplus when the toll is $18


"Consumer surplus = ½ \times base \times height"


"Consumer surplus = ½ \times 100,000 \times (20-18)"


"Consumer surplus =100,000"



C )

Given that

"TC = \$ \space8,000,000"


"Y = 1,000,000 - 50,000P"


 When toll is P = 0


"Y = 1,000,000"



Because when the toll is zero (0) there will be a Consumer surplus of 10,000,000 which is a benefit to the society.